OpenStudy (anonymous):

can someone help me find the solution to this integration: ((5x+8)^1/2+11)/(sqrt5x+8)

OpenStudy (anonymous):

one more time. is this \[\int \frac{\sqrt{5x+8}+11}{\sqrt{5x+8}}dx=\int dx+\int\frac{11}{\sqrt{5x+8}}dx\]

OpenStudy (anonymous):

if so, easy answer is \[x+\frac{22}{5}\sqrt{5x+8}\]

OpenStudy (anonymous):

yes that is the problem

OpenStudy (anonymous):

then that is the answer. easy more or less

OpenStudy (anonymous):

so, if i take the derivative of that answer i should get the orginal problem back right

OpenStudy (anonymous):

comes with my personal guarantee

OpenStudy (anonymous):

, how do i compsate for the chain rule here, and in general with working with integrals

OpenStudy (anonymous):

i am not sure what "compensate for the chain rule" means. wanna take the derivative and see what we get?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the derivative of x is 1. that was easy

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the derivative of \[\sqrt{5x+8}\] is \[\frac{5}{\sqrt{5x+8}}\]

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

the derivative of \[\sqrt{5x+8}\] is \[\frac{5}{2\sqrt{x+8}}\]

OpenStudy (anonymous):

how is that

OpenStudy (anonymous):

so we have \[1+\frac{22}{5}\times \frac{5}{2\sqrt{5x+8}}\]

OpenStudy (anonymous):

the chain rule.

OpenStudy (anonymous):

oh i got it

OpenStudy (anonymous):

how is what? ok the derivative of \[\sqrt{\text{something}}\] is \[\frac{ \text{derivative of something}}{2\sqrt{\text{something}}}\]

OpenStudy (anonymous):

so we have \[1+\frac{11}{\sqrt{5x+8}}\] and i'll be damned if i am doing the algebra, but if you put this over one denominator you will get your original integrand.

OpenStudy (anonymous):

no, i got it

OpenStudy (anonymous):

dont get mad

OpenStudy (anonymous):

not mad at all. you just should have seen the algebra i did earlier

OpenStudy (anonymous):

hours of it because today's problems all required a ton of algebra

OpenStudy (anonymous):

i messed up, cause i didnt break up the term 22sqrt5x+8/5 into 22/5 *sqrt5x+8

OpenStudy (anonymous):

so i just meant i will let you do this one. hope the steps are clear

OpenStudy (anonymous):

ahh.but it is ok yes?

OpenStudy (anonymous):

might be alot to ask but, can youhelp me out mwith one optmization problem

OpenStudy (anonymous):

sure as long as it is over before 4

OpenStudy (anonymous):

judge judy time aka beer o'clock

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

The top and bottom margins of a poster are each 6cm. The side margins are each 4cm. If the printed area on the poster is fixed at 384 cm^2, find the dimensions of the poster with the smallest total area

OpenStudy (anonymous):

total area includes margins right?

OpenStudy (anonymous):

yes, i believe

OpenStudy (anonymous):

ok inside margin call width x and height y, then we know \[xy=384\] making \[y=\frac{384}{x}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

whole area including margin will be \[(x+8)(y+12)\]

OpenStudy (anonymous):

now put \[y=\frac{384}{x}\] to get total area \[A(x)=(x+8)(\frac{384}{x}+12)\]

OpenStudy (anonymous):

multiply this mess out, take derivative set = 0 and solve

OpenStudy (anonymous):

i get \[A(x)=384+12x+\frac{3072}{x}+96\] \[A(x)=12x+\frac{3072}{x}+480\] \[A'(x)=12-\frac{3072}{x^2}\] \[A'(x)=\frac{12x^2-3072}{x^2}\]

OpenStudy (anonymous):

solve \[12x^2-3072=0\] \[12x^2=3072\] \[x^2=256\] \[x=16\] and that is that!

OpenStudy (anonymous):

thanks